OpenStudy (anonymous):

would 4/17 be a proper fraction

OpenStudy (anonymous):

Yes, yes it would.

OpenStudy (kainui):

An improper fraction would be something like 3/2. Three halves is really 1 whole and 1 half, right?

OpenStudy (anonymous):

Right, Kainui. Any time the denominator is greater than the numerator, we have a proper fraction. Otherwise it is improper.

OpenStudy (anonymous):

9. What is the sum of 2/5 and 2/4?

OpenStudy (anonymous):

I don't know how to work these problems can someone explain please

OpenStudy (anonymous):

Consider a common denominator of 4 and 5. A quick way to find a common denominator is to multiply the two denominators together: 4 * 5 = 20. Try rewriting both of the original fractions with a denominator of 20. For example, 2/5 = (2 * 4) / (5 * 4) = 8/20. Try writing the other fraction with a denominator of 20.

OpenStudy (anonymous):

so 8/20 is the answer

OpenStudy (anonymous):

okay i get thanks

OpenStudy (anonymous):

Not quite. 8/20 is the first fraction. Try writing the second fraction with a denominator of 20. Then you can add the two together.

OpenStudy (anonymous):

so 2/5+2/4=8/20+10/20=18/20 is this right and did I do it correctly

OpenStudy (anonymous):

simplify it is 9/10

OpenStudy (anonymous):

Exactly !

OpenStudy (anonymous):

thanks so much